What is Sipho's age and Litha's age in 6 years

1
Expert's answer
2022-03-07T17:25:02-0500

The question is incomplete..It is supposed to be as follows

Sipho and Litha are brothers. The sum of their ages is 29. In 6 years' time, Sipho will be the same age as Litha was 3 years ago. What is Sipho's age and Litha's age in 6 years?


Let X be the age of Sipho and Y to be the age of Litha

Therefore "X+Y=29"

In six years' time

"X+6=Y-3"

"Y=X+9"

Substitute in

"X+Y=29\\X+(X+9)=29\\"

"2X=20\\X=10\\
Y=29-10=19"

Therefore, their current years is;

"sipho=10years\\ litho=19 years"


In six years their age would be:

"sipho=10+ 6 =16 years\\ litho=19 +6 = 25 years" [Answer]